# Variable Substitution and Bilinear Programming for Aligning Partially Overlapping Point Sets

### v1

@@ 8 Experiments
Refer to caption Refer to caption
@@
Figure 6: Average registration errors (top 2 rows) and run times (bottom row) by RPM-BP, RPM-HTB [6], and RPM-CAV [4] under various $n_{p}$ values (ranging from $1/2$ to $/1$ 1 of the ground truth value) over 100 random trials for 2D deformation, noise, mixed outliers and inliers, separate outliers and inliers, and occlusion+outlier tests.
@@
Similar to the experimental setup detailed in Sec. 8.1.1, we conducted five types of tests to assess a method’s resilience against various disturbances. These tests include: i) Deformation test, ii) Noise test, iii) Mixed outliers and inliers test, iv) Separate outliers and inliers test, and v) Occlusion+outlier test. Refer to Fig. 8 for an illustration. Additionally, examples of registration by different methods are depicted in Fig. 10.
The top 2 rows of Fig. 9 displays registration errors from various methods. Results show that RPM-BP and RPM-HTB, with $n_{p}$ chosen as the ground truth, exhibits robustness against deformation, noise, and outliers when outliers are separate from inliers. However, they loses robustness when outliers are mixed with inliers. In contrast, GORE and TEASER++ struggle with handling deformation, noise, and outliers, especially when outliers are mixed with inliers, likely due to their reliance on features. In terms of different choices of $n_{p}$, RPM-BP performs significantly better when $n_{p}$ closely matches the ground truth.
The bottom row of Fig. 9 displays the average run times for various methods. TEASER++ emerges as the fastest, with Go-ICP following closely behind, and RPM-BP and RPM-HTB trailing after. GORE consistently exhibits slower performance across most cases. Notably, FRS experiences a rapid decline in efficiency when faced with challenging registration problems, such as an increase in outliers in both types of outlier tests.
Examining the two types of outlier tests also reveals how different methods scale with problem size. Go-ICP and GORE demonstrate the best scalability with problem size, followed by TEASER++, RPM-BP and RPM-HTB. Conversely, FRS exhibits the poorest scalability with problem size.
Refer to caption Refer to caption
Figure 8: (a) to (c): Model point sets and corresponding examples of scene point sets in the deformation and noise tests, respectively. (d) to (i): Examples of model and scene point sets in the mixed outliers and inliers test ((d), (e)), separate outliers and inliers test ((f), (g)), and occlusion+outlier test ((h), (i)), respectively. In all cases, model points are indicated by red circles, while scene points are represented by blue crosses.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 9: Average registration errors (top 2 rows) and run times (bottom row) of RPM-BP, RPM-HTB, Go-ICP, FRS, GORE, and TEASER++ under varying $n_{p}$ values (ranging from $1/2$ to $/1$ 1 of the ground truth value) across 100 random trials for 3D deformation, noise, mixed outliers and inliers, separate outliers and inliers, and occlusion+outlier tests. For the separate outliers and inliers and occlusion+outlier tests, only a portion of FRS’s run time results is presented, as FRS experiences significant slowdowns with large problem sizes.
@@
Figure 10: Examples of registration results from different methods in the separate outliers and inliers test, where the $n_{p}$ values of RPM-HTB and Go-ICP are both chosen to match the ground truth.
@@
The 3DMatch benchmark [41] encompasses scans of 62 scenes sourced from five established RGB-D reconstruction datasets (sun3d, 7-scenes, rgbd-scenes-v2, bundlefusion, and analysis-by-synthesis). D3Feat [42] offers a unified file structure and format. For convenience, we utilize the scan pairs provided by D3Feat to assess the performance of various methods.
The matching errors from different methods are depicted in Fig. 12. The results indicate that RPM-BP and RPM-HTB generally outperform Go-ICP and FRS but lag behind GORE and TEASER++. The superior performance of GORE and TEASER++ is partly attributed to their use of features, which significantly enhances their capability to match complex structures. In contrast, other methods rely solely on point position information. Additionally, the 3DMatch benchmark comprises purely rigid transformations, aligning well with the assumptions of GORE and TEASER++. Among RPM-BP and RPM-HTB, RPM-BP exhibits slightly superior performance, particularly evident in the sun3d and rgbd-scenes-v2 tests.
Examples of registration by different methods are presented in Fig. 12.
Refer to caption Refer to caption
@@ 9 Conclusion
In this paper, we introduced a BnB-based point set registration algorithm capable of aligning partially overlapping point sets while remaining invariant to the corresponding transformation. The method offers several advantages: the lower bound can be efficiently computed via linear assignment and low-dimensional convex quadratic program, and the dimensionality of the branching space matches the number of transformation parameters. These attributes make the method computationally efficient and scalable with problem size. By employing the BnB algorithm, our method demonstrates robustness against non-rigid deformations, positional noise, and outliers, particularly when outliers are distinct from inliers compared to state-of-the-art approaches.
However, the proposed method is limited to transformations with few parameters due to the exponential complexity of the BnB algorithm with respect to the dimensionality of the branching space. Additionally, the method’s robustness is significantly affected by the parameter setting for $n_{p}$. In the future, we plan to develop an adaptive scheme for setting $n_{p}$.
ACKNOWLEDGMENTS
This work was supported by the Fundamental Research Program of Shanxi Province (202103021223381, 202303021222267 and 202303021222271) and Science and Technological Innovation Program of Higher Education Institutions in Shanxi under Grant (2023L335)
Appendix A A weak result about positive semidefiniteness of matrix $\mathbf{H}^{0}+\mathbf{C}$
@@
Matrix $\mathbf{H}^{0}+\mathbf{C}$ is close to being positive semidefinite.
@@
For any $\mathbf{p}\in\Omega$, it is apparently that $\mathbf{p}\geq 0$. Thus, given $\mathbf{p}\in\Omega$, the function $E$ in equation (3a) is a convex quadratic function of $\boldsymbol{\theta}$. Therefore, from equation (5), we can infer that $\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})+\mathbf{C}\succeq 0$ for any $\mathbf{p}\in\Omega$.
Let
$\mathbf{p}^{ij}=\arg\min_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}$
Apparently ${\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p}^{ij})}+\mathbf{C}\succeq 0$. Based on the fact that $\underline{\Gamma}_{ij}=\min_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}=\min_{\mathbf{p}\in\{\mathbf{p}^{ij}\}}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}$, we have that $\underline{\boldsymbol{\Gamma}}$ is close to one of $\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p}^{ij})$, $\forall i,j$. Consequently $\underline{\boldsymbol{\Gamma}}+\mathbf{C}$ is close to being positive semidefinite.
In an analogous way, let
$\mathbf{q}^{ij}=\arg\max_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}$
Apparently ${\text{mat}(\mathbf{K}\mathbf{B}_{2}q^{ij})}+\mathbf{C}\succeq 0$. Based on $\overline{\Gamma}_{ij}=\max_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}=\max_{\mathbf{p}\in\{\mathbf{q}^{ij}\}}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}$, we have $\underline{\boldsymbol{\Gamma}}$ is close to one of $\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{q}^{ij})$, $\forall i,j$. Consequently $\overline{\boldsymbol{\Gamma}}+\mathbf{C}$ is close to being positive semidefinite.
Therefore $\mathbf{H}^{0}+\mathbf{C}=\frac{\left(\underline{\boldsymbol{\Gamma}}+\mathbf{C}\right)+\left(\overline{\boldsymbol{\Gamma}}+\mathbf{C}\right)}{2}$ is close to being positive semidefinite.

### v2

@@ 5.7 BnB Algorithm Convergence
Figure 4: Upper (first row) and lower bounds (second row) generated in each iteration of RPM-BP. Our method is tested with the $n_{p}$ value chosen as the ground truth and with varying initial ranges of $\boldsymbol{\theta}$: $\boldsymbol{\theta}_{gt}-\Delta,\boldsymbol{\theta}_{gt}+\Delta]$, where $\boldsymbol{\theta}_{gt}$ represents the ground truth $\boldsymbol{\theta}$ solution. Here, the margin $\Delta$ takes values of $0.5$, 11, and $1.5$ respectively.
6 Application 1: 2D Similarity and Affine Registration
Our algorithm is designed to handle various transformations as long as they can be expressed in the form $\mathbf{T}(\mathbf{x}|\boldsymbol{\theta})=\mathbf{J}(\mathbf{x})\boldsymbol{\theta}$. Both 2D similarity and affine transformations fit this structure, allowing for a straightforward application of our framework.
For a 2D similarity transformation, the transformation matrix and parameter vector are:
$\mathbf{T}(\mathbf{x}|{\boldsymbol{\theta}})=\begin{bmatrix}\theta_{1}&-\theta_{2}\\ \theta_{2}&\theta_{1}\end{bmatrix}\begin{bmatrix}x^{1}\\ x^{2} \end{bmatrix}+\begin{bmatrix}\theta_{3}\\ \theta_{4}\end{bmatrix}=\begin{bmatrix}x^{1}&-x^{2}&1&0\\ x^{2}&x^{1}&0&1\end{bmatrix}\boldsymbol{\theta}$  (20)
Here, $\theta_{3},\theta_{4}]^{\top}$ represents the translation, while $\theta_{1}$ and $\theta_{2}$ are derived from the scale $s$ and rotation angle $\phi$ as $\theta_{1}=s\cos(\phi)$ and $\theta_{2}=s\sin(\phi)$. For this transformation, it can be derived that $\mathbf{B}_{2}$ corresponds to rows $[1,3,4]$ of $\mathbf{B}$, and $\mathbf{C}=\text{diag}([0,0,n_{p},n_{p}])$.
A 2D affine transformation is a more general case:
$\mathbf{T}(\mathbf{x}|{\boldsymbol{\theta}})=\begin{bmatrix}\theta_{1}&\theta_{2}\\ \theta_{3}&\theta_{4}\end{bmatrix}\begin{bmatrix}x^{1}\\ x^{2} \end{bmatrix}+\begin{bmatrix}\theta_{5}\\ \theta_{6}\end{bmatrix}=\begin{bmatrix}x^{1}&x^{2}&0&0&1&0\\ 0&0&x^{1}&x^{2}&0&1\end{bmatrix}\boldsymbol{\theta}$  (21)
For this transformation, $\mathbf{B}_{2}$ corresponds to rows $[1,2,5,8,11]$ of $\mathbf{B}$, and $\mathbf{C}=\text{diag}([0,0,0,0,n_{p},n_{p}])$.
7 Application 2: 3D Rigid Registration
Our method is theoretically capable of handling 3D affine transformations, which has the form
[…]
It can be derived that $\mathbf{B}_{2}$ corresponds to rows $[1,2,3,10,14,15,22,27,34]$ of $\mathbf{B}$, and $\mathbf{C}=\text{diag}([\underbrace{0,\ldots,0}_{9\text{ zeros}},n_{p},n_{p},n_{p}])$. Nevertheless, the large number of parameters (12 for a full affine transformation) results in a high-dimensional branching space, significantly slowing down convergence. Therefore, for 3D point cloud registration, we focus on the more constrained rigid transformation.
A 3D rigid transformation involves a rotation and a translation. We represent the 3D rotation using the angle-axis representation, where a rotation is defined by a 3D vector $\mathbf{r}$. The corresponding $3\times 3$ rotation matrix $\mathbf{R}\in\mathbb{SO}_{3}$ can be obtained via the matrix exponential map [21]:
where $\cdot$$_{\times}$ is the skew-symmetric matrix representation:
with $r_{i}$ being the $i$-th element of $\mathbf{r}$.
We adapt our core algorithm from Section 5 to handle this non-linear rigid transformation. The key modifications are:
  1. 1.
    Instead of branching over the full affine parameters, we branch over the 6-dimensional rigid parameters: the angle-axis vector $\mathbf{r}$ and the translation vector $\mathbf{t}$.
  2. 2.
    In each iteration of the BnB algorithm, the range of the rotation matrix $\mathbf{R}$ is computed from the current range of the angle-axis vector $\mathbf{r}$ using the formula in Eq. (22).
For the upper bound calculation, we continue to use the affine-based approach from Eq. (19). While a dedicated method for rigid transformations exists (see Sec. 5.3 of [5]), it often yields higher upper bounds, which hinders the pruning process and slows down our algorithm’s convergence. Therefore, we prioritize speed by using the affine upper bound, even though it doesn’t strictly enforce the rigid constraint.
7.1 Efficiently Computing the Range of $\mathbf{R}$
Given a range $\underline{\mathbf{r}},\overline{\mathbf{r}}]$ for the angle-axis vector $\mathbf{r}$, finding the exact range of each element of the rotation matrix $\mathbf{R}$ is a computationally expensive non-linear optimization problem. Performing this calculation at every step of the BnB algorithm would be a significant bottleneck.
To address this, we use a precomputation and grid-based approximation method. Before the BnB algorithm begins, we create a regular grid over the initial search space of $\mathbf{r}$. For each grid point, we compute the corresponding rotation matrix $\mathbf{R}$ and store the values. Then, during the BnB process, when we need to find the range of $\mathbf{R}$ for a given hypercube $\underline{\mathbf{r}},\overline{\mathbf{r}}]$, we simply find all precomputed grid points that fall within this hypercube and take the minimum and maximum of their corresponding $\mathbf{R}$ values. This provides a fast, though approximate, way to get the necessary bounds, significantly speeding up the overall algorithm.
8 Experiments
We implemented the proposed algorithm, which we refer to as RPM-BP, using Matlab R2023b. All comparative experiments were conducted on a computer equipped with 6-core CPUs running at 3.2 GHz.
For competing methods that output only point correspondences, we used these correspondences to compute the optimal affine transformation between the two point sets. We defined the matching error as the root-mean-square distance between the transformed model inliers and their corresponding scene inliers.
To ensure efficient computation, the maximum branching depth of RPM-BP was set to 12.
8.1 2D Registration
We compared our method with RPM-HTB [5], RPM-PA [4], and RPM-CAV [3]. These methods were chosen because they are all based on global optimization techniques, can handle partial overlap, and allow for arbitrary similarity transformations, making them suitable for direct comparison with our approach.
8.1.1 2D Synthetic Data
Synthetic data allows us to systematically evaluate an algorithm’s resilience to specific disturbances. We performed five distinct types of tests:
  1. 1.
    Deformation test: The prototype shape is non-rigidly deformed to create the scene point set.
  2. 2.
    Noise test: Positional noise is added to the prototype shape to generate the scene point set.
  3. 3.
    Mixed outliers test: Random outliers are superimposed on both the prototype and scene shapes.
  4. 4.
    Separate outliers test: Random outliers are added to different areas of the prototype shape to generate the two point sets.
  5. 5.
    Occlusion + Outlier test: The prototype shape is occluded to produce two point sets, to which random outliers (with a fixed outlier-to-data ratio of 0.5) are then added in different regions.
These tests are visually illustrated in Fig. 7. To assess a method’s ability to handle arbitrary rotations and scaling, random rotations and scalings within the range of $[0.5,1.5]$ were applied when generating the point sets. Examples of registration results from different methods are provided in Fig. 7.
The registration errors produced by each method are shown in the top two rows of Fig. 7. Our results indicate that RPM-BP and RPM-HTB, when the number of correspondences $n_{p}$ is chosen as the ground truth, are robust against deformation, noise, and separate outliers. However, their performance degrades when outliers are mixed with inliers. This limitation is a consequence of our method not being $\epsilon$-globally optimal, as discussed in Sec. 5.7.
When using different transformation types, RPM-BP demonstrates greater resilience to disturbances when employing a similarity transformation compared to an affine transformation. Furthermore, the performance of RPM-BP is significantly influenced by the chosen $n_{p}$ value, with results improving as $n_{p}$ approaches the ground truth. We also found that RPM-BP is less sensitive to variations in $n_{p}$ than RPM-HTB, especially in outlier and occlusion tests.
The average run times are depicted in the bottom row of Fig. 7. RPM-BP is the most efficient method in the deformation, positional noise, and mixed inlier-outlier tests. Comparing transformation types, RPM-BP is more efficient with similarity transformations due to the smaller number of parameters.
Finally, the outlier tests also provide insight into how each method scales with problem size. RPM-BP, RPM-HTB, and RPM-CAV demonstrate the best scalability, followed by RPM-PA.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ 8 Experiments
Figure 6: Average registration errors (top 2 rows) and run times (bottom row) by various methods under various $n_{p}$ values (ranging from $1/2$ to $1/1$ of the ground truth value) over 100 random trials for 2D deformation, positional noise, mixed outliers and inliers, separate outliers and inliers, and occlusion+outlier tests.
@@
8.1.2 2D Real-World Data
Point sets extracted from images represent a realistic and challenging data type for evaluating the robustness and generalizability of registration methods, as they inherently contain noise, irregular contours, and complex structural features.
Experimental Setup
To generate these point sets, we utilized the standard Canny edge detector on selected images sourced from the Caltech-256 [32] and VOC2007 [33] datasets. A critical test of rotational invariance was implemented by subjecting the model point set to a deliberate $180^{\circ}$ rotation prior to registration against the scene point set. For consistency, a Similarity transformation model was assumed and used for all tested methods. The visual setup for this experiment is detailed in Fig. 9, with final alignment results shown in Fig. 9.
Results
The registration outcomes presented in Fig. 9 demonstrate that our method consistently achieves a tighter alignment to the underlying target objects than competing algorithms. This advantage is notably pronounced in the bike image test.
Figure 8: Model images with model point sets superimposed.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 9: For each category: scene images with scene point sets superimposed, registration results by RPM-BP, RPM-HTB [5], RPM-PA [4] and RPM-CAV [3] using similarity transformation. The $n_{p}$ value for each method is chosen as $0.9$ the minimum of the cardinalities of two point sets.
@@ Results
Similar to the 2D experiments, we conducted five types of tests to evaluate each method’s resilience to various disturbances: i) Deformation, ii) Noise, iii) Mixed outliers, iv) Separate outliers, and v) Occlusion + Outlier tests. Fig. 12 provides a visual illustration of these tests, while Fig. 12 shows examples of registration results from different methods.
The registration errors for various methods are shown in the top two rows of Fig. 12. The results indicate that RPM-BP and RPM-HTB, when $n_{p}$ is set to the ground truth value, are robust against deformation, noise, and separate outliers. However, their robustness diminishes when outliers are mixed with inliers. In contrast, GORE and TEASER++ struggle with deformation, noise, and mixed outliers, likely because they rely on features. As in the 2D case, the performance of RPM-BP improves significantly when $n_{p}$ closely matches the ground truth.
The bottom row of Fig. 12 shows the average run times. TEASER++ is the fastest, followed by Go-ICP, and then RPM-BP and RPM-HTB. GORE is consistently slower in most cases, while FRS experiences a rapid decrease in efficiency when the registration problem becomes more challenging, such as with an increasing number of outliers.
In terms of scalability with problem size, Go-ICP and GORE perform best, followed by TEASER++, RPM-BP, and RPM-HTB. FRS, on the other hand, shows the poorest scalability.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: (a) to (c): Model point sets and examples of scene point sets in the deformation and noise tests, respectively. (d) to (i): Examples of model and scene point sets in the mixed outliers and inliers test ((d), (e)), separate outliers and inliers test ((f), (g)), and occlusion+outlier test ((h), (i)), respectively. In all cases, model points are indicated by red circles, while scene points are represented by blue crosses.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Average registration errors (top 2 rows) and run times (bottom row) of various methods under varying $n_{p}$ values ($1/2$ or $1/1$ of the ground truth value) across 100 random trials for 3D deformation, positional noise, mixed outliers and inliers, separate outliers and inliers, and occlusion+outlier tests.
@@
Figure 12: Examples of registration results from different methods in the separate outliers and inliers test, where the $n_{p}$ values of RPM-HTB and Go-ICP are both chosen as the ground truth.
@@
The 3DMatch benchmark [34] provides scans from 62 scenes across five established RGB-D reconstruction datasets (sun3d, 7-scenes, rgbd-scenes-v2, bundlefusion, and analysis-by-synthesis). For convenience, we used the scan pairs from D3Feat [35] to evaluate the performance of our method.
Fig. 14 shows the matching errors for the different methods. The results indicate that RPM-BP and RPM-HTB generally outperform Go-ICP and FRS but are less accurate than GORE and TEASER++. The superior performance of GORE and TEASER++ can be attributed to their use of features, which significantly enhances their ability to match complex structures. In contrast, the other methods rely solely on point position information. Furthermore, the 3DMatch benchmark contains purely rigid transformations, which aligns perfectly with the assumptions of GORE and TEASER++. Between our method and RPM-HTB, RPM-BP demonstrates slightly better performance, especially in the sun3d and rgbd-scenes-v2 tests.
Examples of registration results are presented in Fig. 14.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ 9 Conclusion
In this paper, we presented RPM-BP, a novel Branch-and-Bound (BnB) algorithm for globally optimal point set registration under partial overlap. Our method is designed to be invariant to the underlying transformation, encompassing both 2D similarity/affine and 3D rigid transformations.
The efficiency of RPM-BP stems from several key features:
  1. 1.
    Efficient Lower Bound Computation: The lower bound is efficiently computed by decoupling the minimization problem into a tractable linear assignment problem (for correspondences) and a low-dimensional convex quadratic program (for transformation parameters).
  2. 2.
    Low-Dimensional Branching Space: The dimensionality of the branching space is confined solely to the transformation parameters, which is inherently low-dimensional (e.g., 4 for 2D similarity, 6 for 3D rigid). This greatly mitigates the exponential complexity typical of BnB algorithms.
These attributes contribute to the method’s computational efficiency and its strong scalability with respect to the number of points. Our extensive experiments confirm that RPM-BP demonstrates superior robustness against non-rigid deformations, positional noise, and outliers, particularly in scenarios where outliers and inliers are spatially separated.
Despite its advantages, the proposed method has two primary limitations. First, its use is inherently restricted to transformations with a small number of parameters, as the complexity of the BnB algorithm remains exponential with respect to the branching space dimensionality. Second, the method’s overall robustness is sensitive to the parameter setting for the number of inliers, $n_{p}$.
For future work, we plan to focus on developing an adaptive scheme for setting the parameter $n_{p}$ to enhance the method’s performance and practical applicability.
Appendix A Heuristic Justification for the Positive Semidefiniteness of $\mathbf{H}^{0}+\mathbf{C}$
The convergence of our Branch-and-Bound algorithm relies on the lower bound subproblem being convex, which requires the matrix $\mathbf{H}^{0}+\mathbf{C}$ to be Positive Semidefinite (PSD). While a formal proof that $\mathbf{H}^{0}+\mathbf{C}\succeq 0$ is elusive, we offer a heuristic justification, consistent with our empirical findings in Section 5.6.
@@ Appendix A Heuristic Justification for the Positive Semidefiniteness of $\mathbf{H}^{0}+\mathbf{C}$
The matrix $\mathbf{H}^{0}+\mathbf{C}$ is an average of matrices that are heuristically expected to be close to Positive Semidefinite.
@@
We first recall the definition of the Hessian matrix $\mathbf{G}(\mathbf{p})$ with respect to $\boldsymbol{\theta}$ for a fixed correspondence vector $\mathbf{p}$:
$\mathbf{G}(\mathbf{p})=\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})+\mathbf{C}.$
Since the point correspondence vector $\mathbf{p}$ satisfies $\mathbf{p}\geq 0$ (i.e., $\mathbf{p}\in\Omega$), and the original objective function $E(\mathbf{p},\boldsymbol{\theta})$ represents a weighted sum of squared errors, it is fundamentally a convex quadratic function of $\boldsymbol{\theta}$ for any feasible $\mathbf{p}$. This mandates that the Hessian must be Positive Semidefinite for all valid $\mathbf{p}$:
$\mathbf{G}(\mathbf{p})\succeq 0\quad\text{for all }\mathbf{p}\in\Omega.$
The matrix $\mathbf{H}^{0}+\mathbf{C}$ is defined as the average of two bounding matrices:
$\mathbf{H}^{0}+\mathbf{C}=\frac{1}{2}\left[(\underline{\boldsymbol{\Gamma}}+\mathbf{C})+(\overline{\boldsymbol{\Gamma}}+\mathbf{C})\right].$
where $\underline{\boldsymbol{\Gamma}}+\mathbf{C}$ (and similarly $\overline{\boldsymbol{\Gamma}}+\mathbf{C}$) is an element-wise lower bound on the Hessian $\mathbf{G}(\mathbf{p})$:
$(\underline{\boldsymbol{\Gamma}}+\mathbf{C})_{ij}=\left(\min_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}\right)+\mathbf{C}_{ij}.$
The element-wise minimum $\underline{\boldsymbol{\Gamma}}_{ij}$ (and maximum $\overline{\boldsymbol{\Gamma}}_{ij}$) is attained by some $\mathbf{p}\in\Omega$. Since all $\mathbf{G}(\mathbf{p})$ are PSD matrices, the elements used to construct $\underline{\boldsymbol{\Gamma}}+\mathbf{C}$ and $\overline{\boldsymbol{\Gamma}}+\mathbf{C}$ are drawn from a collection of PSD matrices. While $\underline{\boldsymbol{\Gamma}}+\mathbf{C}$ and $\overline{\boldsymbol{\Gamma}}+\mathbf{C}$ may not be strictly PSD themselves (as the minimum/maximum of each entry may be attained by a different $\mathbf{p}$), they are structurally derived from PSD matrices.
It is therefore reasonable to conclude that both $(\underline{\boldsymbol{\Gamma}}+\mathbf{C})$ and $(\overline{\boldsymbol{\Gamma}}+\mathbf{C})$ are close to being positive semidefinite. Because the sum of two PSD matrices is PSD, their average, $\mathbf{H}^{0}+\mathbf{C}$, is consequently expected to be close to being positive semidefinite, ensuring the practical convexity of the subproblem. ∎